### v1

# Early Explorations of Recommender Systems for Physical Activity and Well-being

Abstract
As recommender systems increasingly guide physical actions, often through wearables and coaching tools, new challenges arise around how users interpret, trust, and respond to this advice. This paper introduces a conceptual framework for tangible recommendations that influence users’ bodies, routines, and well-being. We describe three design dimensions: trust and interpretation, intent alignment, and consequence awareness. These highlight key limitations in applying conventional recommender logic to embodied settings. Through examples and design reflections, we outline how future systems can support long-term well-being, behavioral alignment, and socially responsible personalization.
1 Introduction and Background
Recommender systems have long shaped how users engage with digital media, products, and information. They are now embedded in wearables, health apps, and coaching tools that influence how we move, rest, and manage our bodies. In these settings, recommendations are not just digital suggestions. They prompt physical actions, such as what to eat, when to train, or how far to run. We refer to these as tangible recommendations, where the consequences are embodied, cumulative, and often invisible to the algorithms that produce them.
As these systems guide offline behavior, new questions arise. What does it mean to trust a system that recommends physical activity? How can intent be inferred or adjusted when user goals are fluid and context dependent? How should systems account for the longer-term consequences of their suggestions?
These questions are especially relevant in domains like health, sport, and well-being, where motivation, capacity, and priorities shift frequently. A system that nudges exertion may improve fitness in one case and cause harm in another. Yet most recommender systems remain focused on digital engagement, often optimizing for click-through or novelty rather than alignment with bodily needs.
The field is beginning to address concerns around fairness, explainability, and social impact. Several recent efforts call for evaluation frameworks that move beyond accuracy and toward broader public interest goals [6]. Work on health-aware systems [10, 3] and behaviorally informed recommendation [4] points in this direction. Still, we lack conceptual tools to understand how users interpret and act on recommendations in physical contexts.
This position paper addresses the need for recommender systems that account for the physical, emotional, and social dimensions of user experience. We focus on three persistent challenges in this space: trust, intent, and consequence. These themes are explored through a conceptual lens and grounded in well-being-oriented use cases. Our goal is to support the design of recommender systems that align with embodied needs, promote long-term well-being, and enable responsible personalization.
2 Framework: Dimensions of Tangible Recommendation
Traditional recommender systems measure utility through relevance, engagement, or predictive accuracy. When systems begin to influence physical behavior, such as when and how users exercise, these metrics fall short. Tangible recommendations raise distinct design concerns. Building on prior work in health-aware RS, HCI, and AI ethics, we propose a framework comprising three dimensions that define the design space of tangible recommendation: trust and interpretation, intent and alignment, and consequence awareness. Each dimension surfaces key tensions in designing responsible systems for physical activity and well-being, as outlined below and shown in Table˜1.
2.1 Trust and Interpretation
Trust is essential for users to act on system suggestions, especially when those suggestions affect the body. Trust is not static; it evolves through interaction, interpretation, and experience [8]. In tangible settings, trust depends on how users perceive system intent, how well the recommendation aligns with their state or knowledge, and how transparently it is framed.
Work in explainable AI and recommender systems shows that explanation style strongly influences trust and compliance [7]. Designing for interpretive trust involves communicating uncertainty, providing rationale, and acknowledging limits. A confident but inappropriate recommendation, e.g., such as prompting exertion during fatigue, can damage trust more than a cautious, qualified one. Presenting suggestions as guidance rather than prescriptions supports longer-term engagement [9].
2.2 Intent and Alignment
Recommendations presume an underlying user goal, such as improving health, gaining strength, or recovering from injury. In embodied settings, intent is often fluid, context sensitive, and difficult to infer. People regularly shift between goals based on mood, fatigue, or external events. Misalignment between inferred and actual intent can lead to frustration or harm [5].
To address this, systems must support changing intent. This could involve explicit tagging (e.g., “today I want to recover”), passive signals (e.g., reduced heart rate variability), or hybrid models. Without such flexibility, recommendations risk becoming irrelevant or counterproductive.
2.3 Consequence Awareness
Tangible recommendations have effects that go beyond engagement. Physical outcomes (e.g., injury), emotional responses (e.g., guilt), and social dynamics (e.g., peer pressure) are rarely captured in standard evaluation metrics [1].
Sustainable systems must be sensitive to downstream effects, even if they cannot predict them exactly. Prior work in health-aware RS emphasizes the need to balance engagement with long-term well-being [10], and recent evaluation research highlights the importance of accounting for context-specific harms [2]. This could include support for recovery, more varied recommendations, and mechanisms that encourage reflection [1, 10, 2].
Table 1: Design dimensions for tangible recommendations
Dimension Design Consideration
Trust and Interpretation Support transparency, calibrate confidence, communicate uncertainty; frame recommendations as guidance, not prescription.
Intent and Alignment Enable dynamic goal-setting and intent signaling; design for fluid, situational motivation rather than static profiles.
Consequence Awareness Consider physical and emotional impact; support recovery, routine variation, and reflective engagement with long-term outcomes.
3 Illustrative Scenarios
To ground the framework, we present two scenarios that illustrate how tangible recommendations relate to trust, intent, and consequence in everyday physical activity. While hypothetical, they reflect patterns from real-world use.
3.1 Scenario 1: Overriding the Body
Maria, a recreational runner, uses a smartwatch with a built-in virtual coach. After a week of intense training, the system recommends a high-intensity interval workout based on her heart rate variability data. Although Maria feels physically exhausted, she interprets the recommendation as a sign she should push harder. She completes the session but suffers muscle strain that prevents her from running for two weeks.
This scenario highlights the risks of over-trusting system authority and the absence of bodily context in the recommendation logic. It exemplifies failures in both interpretive trust and consequence awareness [8, 10].
3.2 Scenario 2: Misaligned Recovery Intent
Carl is recovering from a knee injury. His wearable has learned his preference for long-distance cycling and continues to recommend hilly rides. However, his physiotherapist advises low-resistance indoor training. Without a way to express this change in intent, Carl disregards the system altogether.
This case shows how static user modeling fails with shifting needs. The lack of support for goal fluidity and intent realignment leads to disuse and distrust [5].
3.3 Design Insights
These scenarios show how tangible recommendations challenge assumptions about personalization and system authority. In Maria’s case, trust becomes problematic not due to system failure, but because persuasive recommendations override her bodily awareness. The system is technically correct but contextually misaligned. In Carl’s case, the system misses a shift in intent, leading to irrelevant advice and eventual disengagement. Both examples point to the need for calibrated trust. Trust depends not just on accuracy, but on how well users feel the system reflects their lived context.
Intent must also be modeled as fluid. Physical condition, motivation, and external constraints often shift in ways that static personalization does not capture. Without mechanisms for users to revise or suppress inferred goals, recommendations may become counterproductive. These scenarios also highlight that the consequences of recommendation are not only behavioral, but also emotional and social. When systems ignore recovery needs, reinforce pressure, or limit autonomy, they may erode wellbeing rather than support it.
Tangible recommender systems should be designed with contextual awareness, flexible intent modeling, and sensitivity to consequence. Supporting wellbeing in physical activity involves more than matching preferences. It requires alignment with users’ changing capacities, goals, and interpretations of movement.
4 Positioning and Outlook
The scenarios discussed above demonstrate how recommender systems in embodied contexts raise new challenges for personalization. Although wearables increasingly offer activity suggestions, we still know little about how different types of users interpret and act on this guidance. Athletes, patients, and casual exercisers may have very different relationships to system authority, bodily feedback, and recovery needs. Addressing this gap requires empirical approaches that attend to lived experience, including interviews, diary methods, and reflective data collection. System design must also evolve to treat intent as dynamic. Users should be able to shift, suspend, or redefine their goals in ways that reflect current physical and emotional states.
Evaluating these systems will require a rethinking of priorities. Metrics such as click-through rate or precision offer little insight into the value or risk of a recommendation that affects the body. Future work should explore evaluation approaches that reflect long-term wellbeing, behavioral alignment, and user agency. These include consequence-aware evaluation techniques, subjective trust assessments, and transparency methods that help users understand what is being recommended and why.
This framework contributes to a broader research agenda in sustainable and responsible recommendation. In the domain of physical activity, sustainability refers not to computational efficiency, but to the preservation of health, motivation, and trust over time. Designing for these outcomes means attending to users’ fluctuating goals, embodied constraints, and personal meanings of movement.
By introducing the dimensions of trust, intent, and consequence, this paper outlines a conceptual framework for analyzing and designing recommender systems that guide physical action. Future work will expand on this foundation through empirical study and design. The goal is to support technologies that are both adaptive and respectful of users’ lived realities.